how would you prepare 500ml of 300ppm Cl- solution starting with sodium chloride.

plase elaborate each steps
Answer #1

Concentration of Cl- ion = 300 ppm

1ppm means 1 mg per 1litre

1ppm = mg/L

1ppm =10^-3 g/L

molar mass of NaCl = 58.44 gram/mole

300 ppm = 300 x10^-3 g/L = 300 x10^-3/58.44 mol/L

300ppm = 0.0051 mol/L

concentration of solution = 0.0051 mol/L = 0.0051M

volume of the soluion = 500 ml = 0.500L

Concentration of the solution = number of moles /volume in L

number of moles = concentration x volume = 0.0051 x 0.500 = 0.00255 moles

number of moles of Cl- ion = 0.00255 moles

molar mass of Ci- ion = 35.45 gram/mole

mass of one mole of Cl- = 35.45 grams

mass of 0.00255 moles of Cl- = 0.00255 x 35.45 = 0.0904 grams

mass of Cl- ion = 0.0904 grams.
Answer #2

To prepare a 500 mL solution of 300 ppm Cl⁻ using sodium chloride (NaCl), follow these steps:


Step 1: Understand the definition of ppm.

  • 300 ppm means 300 mg of Cl⁻ per liter of solution.

Mass of Cl⁻ required for 500 mL:\text{Mass of Cl⁻ required for 500 mL:}Mass of Cl⁻=300mg/L×5001000=150mg=0.150g\text{Mass of Cl⁻} = 300 \, \text{mg/L} \times \frac{500}{1000} = 150 \, \text{mg} = 0.150 \, \text{g}


Step 2: Calculate the required amount of NaCl.

  • Molar mass of NaCl:

Molar mass of NaCl=58.44g/mol\text{Molar mass of NaCl} = 58.44 \, \text{g/mol}

  • Molar mass of Cl⁻:

Molar mass of Cl⁻=35.45g/mol\text{Molar mass of Cl⁻} = 35.45 \, \text{g/mol}

To calculate how much NaCl contains 0.150 g of Cl⁻:

Mass of NaCl required:58.4435.45×0.150g=58.44×0.15035.45\text{Mass of NaCl required:} \quad \frac{58.44}{35.45} \times 0.150 \, \text{g} = \frac{58.44 \times 0.150}{35.45}Mass of NaCl0.247g\text{Mass of NaCl} \approx 0.247 \, \text{g}


Step 3: Prepare the solution.

  1. Weigh 0.247 g of NaCl accurately using a balance.

  2. Transfer the NaCl to a 500 mL volumetric flask.

  3. Add distilled water to dissolve the NaCl.

  4. Fill the flask to the 500 mL mark with distilled water.

  5. Mix thoroughly to ensure the salt is completely dissolved.

  •  You now have a 500 mL solution of 300 ppm Cl⁻.


answered by: Monu Kumar Gupta